# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis

### v3

@@ 4.2 Revisiting $\mathcal{KL}$ functions : the class $\mathcal{KL}_{\rm gen}$
We extend the class of $\mathcal{K}\mathcal{L}$ functions which classically serve for dynamical systems stability. In our context, we have a similar use and our $\mathcal{KL}_{\rm gen}$ class constitutes the main part of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds of $\varphi(T^{k}(x))$. We prove that such upper bounds exist if and only if $\nu\in\Lambda$. Moreover, if ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, we are able to construct a function $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$ and such that $e^{-1}\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$.
@@
Unfortunately, in practice, to prove that a pair $(\gamma,\theta)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ seems to be as difficult as proving that a pair $(h,\beta)$ belongs to $\mathbb{\Gamma}(\nu)$. We thus propose a new type of Lyapunov functions for which the conditions does not involve the whole trajectories of the analyzed dynamical system.
@@ 5 Lyapunov type Functions for Discrete-time Systems Maximization
Classically, Lyapunov functions are used as certificates of stability (see  [Ela05] or [KP01]). They bring a constructive approach to prove the stability of the dynamical system through converse Lyapunov theorems which prove that stability notions are equivalent to the existence of Lyapunov functions [Kha02]. For some linear, polynomial or piecewise polynomial systems, Lyapunov functions can be computed from numerical optimization solvers based on linear or semidefinite programming (for example see [GH15] and references therein).
Classical quadratic Lyapunov functions solve our problem when the dynamical system is stable and affine and the objective function is quadratic [Adj21]. In Subsection 5.1, we generalize this approach theoretically. Then in Subsection 5.2, we define a more general class of Lyapunov functions. The existence of those functions does not ensure the stability of the system but just the finiteness of the limsup of the sequence $\nu$ defined at Eq. (1). Last but not least, we etablish an equivalence theorem between the existence of those new Lyapunov functions and the existence of an element $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
5.1 Classical Lyapunov Functions Theory
5.1.1 Definition and Useful Facts
[…]
Definition 7 (Classical Lyapunov Functions).
A function $V:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is said to a Lyapunov function if and only if :
  1. 1.
    There exist $\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ such that:
    $\alpha_{1}\circ\|\cdot\|\leq V\leq\alpha_{2}\circ\|\cdot\|$
  2. 2.
    There exists $\lambda\in(0,1)$ such that
    $V\circ T\leq\lambda V$
[…]
  1. 1.
    there exists $\alpha:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ positive definite111a function $g:X\mapsto\mathbb{R}_{+}$ where $X\subset\mathbb{R}^{d}$ and $0\in X$ is said to be positive definite (on $X$) if $g(0)=0$ and $g(x)>0$ for all $x\in X\backslash\{0\}$ and continuous such that $V\circ T\leq V-\alpha\circ\|\cdot\|$;
  2. 2.
    there exists $\rho:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ positive definite, continuous and such that $\mathrm{Id}-\rho$ is positive definite satisfying $V\circ T\leq\rho\circ V$.
We will propose a proof for this equivalence later on the new class of Lyapunov functions used for the maximization problem.
First, we remark that from the first statement of Def. 7, $V$ is positive definite. If we require $V$ to be continuous and such that $\lim V(x)=+\infty$ as $\|x\|\to+\infty$ (see e.g. [Adj24]), then the first statement of Def. 7 is actually equivalent to $V$ is postive definite. Second, if there exists a classical Lyapunov function for $T$, we $T$ must be null at zero [^2] Actually, if $T(e)=e$ for a nonzero vector, we could replace $V$ by $V(\cdot-e)$ and the lower and the upper bounds on $V(\cdot-e)$ by respectively $\alpha_{1}(\|\cdot-e\|)$ and $\alpha_{2}(\|\cdot-e\|)$. Finally, from the second statement of Def 7, we can define a kind of norm operator relative to a Lyapunov function. We have, for all Lyapunov function $V$ associated to the dynamic $T$, for all $k\in\mathbb{N}$:
$N_{T^{k}}(V):=\sup_{\begin{subarray}{c}x\in\mathbb{R}^{d}\\ x\neq 0\end{subarray}}\dfrac{V(T^{k}(x))}{V(x)}\in(0,1)$ (5)
We will simply write $N_{T}(V)$ when $k=1$. By definition, we have for all $x\in\mathbb{R}^{d}$, $V(T(x))\leq N_{T}(V)V(x)$. Using this and the recurrence principle, we get, for all $k\in\mathbb{N}$, $N_{T^{k}}(V)\leq(N_{T}(V))^{k}$.
5.1.2 A Direct Classical Lyapunov Theorem
Let $f$ be a function from $\mathbb{R}^{d}$ to $\mathbb{R}$. We recall that the function from $\mathbb{R}_{+}$ to $\mathbb{R}$:
$\alpha^{f}:s\mapsto\sup_{\|y\|\leq s}f(y)$ (6)
is continuous when $f$ is continuous (see e.g. [Kha02] or extensions in [Adj24]). Obviously, the function $\alpha^{f}$ is increasing. Moreover, if $f(0)=0$ then $\alpha^{f}$ is nonnegative and satisfies $\alpha^{f}(0)=0$. Last but not least, we have for all $x\in\mathbb{R}^{d}$, $f(x)\leq\alpha^{f}(\|x\|)$.
Theorem 7 (Classical Lyapunov to $\mathbb{\Gamma}(\nu)$).
Assume that $\varphi$ is continuous and there exists a Lyapunov function for $T$. Then there exists $(h,\beta)\in\mathbb{\Gamma}(\nu)$ such that $h(0)=0$. Thus, if Assumption 2 holds, $h$ is useful for $\nu$.
Proof.
If there exists a Lyapunov function $V$ for $T$, there exists $\alpha_{1}\in\mathcal{K}_{\infty}$ such that $\alpha_{1}\circ\|\cdot\|\leq V$. Recall that we have assumed that $\varphi(0)=0$. With the continuity assumption of $\varphi$, we conclude that $\alpha^{\varphi}$ (defined at Eq. (6)) is nonnegative, increasing, continuous and is equal to 0 at 0. Two small issues : $\alpha^{\varphi}$ is not guaranteed to be strictly increasing and surjective from $\mathbb{R}_{+}$ to $\mathbb{R}_{+}$ (the limit of $\alpha^{\varphi}$ at +$\infty$+ ∞ can be finite). We thus add to $\alpha^{\varphi}$ a function $\psi\in\mathcal{K}_{\infty}$ and we have for $y\in\mathbb{R}^{d}$, $\varphi(y)\leq(\alpha^{\varphi}+\psi)(\|y\|)$ and $\alpha:=\alpha^{\varphi}+\psi$ belongs to $\mathcal{K}_{\infty}$.
From $\alpha_{1}\circ\|\cdot\|\leq V$ and $\alpha_{1}\in\mathcal{K}_{\infty}$, we get $\varphi\leq\alpha\circ\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}\circ V$ as $\alpha$ is increasing. Moreover, we have, for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $V(T^{k}(x))\leq N_{T^{k}}(V)V(x)\leq(N_{T}(V))^{k}V(x)\leq(N_{T}(V))^{k}\overline{V_{X^{\mathrm{in}}}}$. As $\alpha$ and ${{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}$ are increasing, we have for all $k\in\mathbb{N}$, $\nu_{k}\leq\alpha\left(\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}\left(N_{T}(V))^{k}\overline{V_{X^{\mathrm{in}}}}\right)\right)$. We thus define $h:x\mapsto\alpha\left(\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}(x\overline{V_{X^{\mathrm{in}}}})\right)$ and $\beta=N_{T}(V)$. We have $h(0)=0$ and we conclude that $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$ when Assumption 2 holds. ∎
The goal of extending the concept of Lyapunov functions is to weaken the hypothesis of Th 7 by removing the continuity and the positivity assumptions on $\varphi$. The extension also allows to prove the equivalence between the existence of useful pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ and the existence of this new type of Lyapunov functions.
5.2 Opt-Lyapunov Functions for Discrete-time Systems Maximization
@@ 5.2.1 Definitions and Useful Facts
A function $V$ is said to be an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ if and only if:
@@
Example 5.
Let us consider the dynamical system of the example of Subsection 2.2. Recall that $\mu\in(0,1/3)$ and the set of initial conditions is :
$X^{\mathrm{in}}=([2\mu,1]\times[\mu,1])\cap\operatorname{Vect}\{(2,1)^{\intercal}\}$
and the dynamics is the following:
$\mathbb{R}^{2}\ni(x,y)\mapsto T(x,y)=A\begin{pmatrix}x\\ y\end{pmatrix}\text{ where }A=\begin{pmatrix}1&1\\ \frac{1}{4}&1\end{pmatrix}$
The function defined, for all $x=(x_{1},x_{2})^{\intercal}\in\mathbb{R}^{2}$ as follows:
$V(x)=\left\{\begin{array}[]{cr}\dfrac{2\mu^{2}}{x_{1}x_{2}}&\text{ if }x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}\\ \\ 0&\text{ otherwise}\end{array}\right.$
is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
Let us prove that it is true. Let $x\in\mathbb{R}^{2}$. If $x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$, then $x=(2,1)^{\intercal}y$ with $y\in\mathbb{R}^{*}$. First, the coordinates of $x$ has the same sign and so $V$ is nonnegative. The fact that $V(x)$ is nonnegative when for $x\notin\operatorname{Vect}\{(2,1)^{\intercal}\}$ or $x=0$ is obvious. Second, let $x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$ and let $y\neq 0$ such that $x=(2,1)^{\intercal}y$. Then, $V(T(x))=V((3y,3y/2)^{\intercal})=(4\mu^{2})/(9y^{2})$ and $V(x)=\mu^{2}/y^{2}$ then $V(T(x))/V(x)=(4y^{2})/(9y^{2})=4/9$. Any value in $[4/9,1)$ can be chosen. Let us take $5/9$. As we have $T(z)\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$ if and only if $z\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$, we conclude that for all $z\notin\operatorname{Vect}\{(2,1)^{\intercal}\}$ or $z=0$, we have $V(T(z))=V(z)=0$. We conclude that $V(T(z))\leq 5V(z)/9$ for all $z\in\mathbb{R}^{2}$. Finally, since $\mu>0$, we have $\max_{x\in X^{\mathrm{in}}}V(x)=\max_{y\in[\mu,1/2]}(2\mu^{2})/(2y^{2})^{-1}=(2\mu^{2})(2\mu^{2})^{-1}=1$.
@@
If $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ then :
@@
2. Let $x$ be a fixed point of $T$. If $V(x)=+\infty$, there is nothing to prove. Now, suppose that $V(x)$ is finite. From the third statement, $0\leq V(T(x))=V(x)\leq\lambda V(x)$ for some $\lambda\in(0,1)$. Hence, $V(x)(1-\lambda)=0$ and we conclude that $V(x)=0$. ∎
@@
If $X^{\mathrm{in}}$ is included in the fixed point set of $T$ then a function $V:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that $V\circ T\leq\lambda V$ for all some $\lambda\in(0,1)$ cannot satisfy $\overline{V_{X^{\mathrm{in}}}}>0$.
Corollary 4 extends the fact that if a classical Lyapunov function (as defined in Def 7) exists for a self-map $T$, $T$ cannot have non-zero fixed point. If an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ exists, $X^{\mathrm{in}}$ contains a vector $x$ such that $x\neq f(x)$.
Prop. 15 shows that the set of Opt-Lyapunov functions contains the classical Lyapunov functions.
Proposition 15 (Classical is Opt).
If $V$ is a classical Lyapunov function for $T$ and $X^{\mathrm{in}}$ is not reduced to $\{0\}$, then $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
Actually, we can replace equivalently the second statement of Def. 8 by $\overline{V_{X^{\mathrm{in}}}}$ is finite and strictly positive. Indeed, with this condition, we recover an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ with $\overline{W_{X^{\mathrm{in}}}}\in(0,1]$ by taking $W:=V/\overline{V}$.
@@
The existence of a function $W:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that :
@@
is equivalent to the existence of a Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
@@
Lemma 4 is a slighty modified version of [Kel14, Lemma 25] for which $f-\mathrm{Id}$ is supposed to be in $\mathcal{K}_{\infty}$. To be self-contained, we propose in Appendix a proof inspired by the one provided in [Kel14].
@@
Again a proof of Lemma 5 is given in [Kel14]. A proof is provided in Appendix for the convenience of the reader.
Proof of Prop. 16.
Suppose that such $W$ and $\rho$ exist. Let $\lambda\in(0,1)$. From Lemma 5, there exists $g$ such that $g(\rho(x))=\lambda g(x)$ for all $x\in\mathbb{R}_{+}$. By extension, we define $g(+\infty)=+\infty$. Then as $g$ is increasing, we have $g(W(T(x)))\leq g(\rho(W(x))=\lambda g(W(x))$ for all $x\in\mathbb{R}^{d}$. Now, as $g$ is strictly increasing there exists some $x\in X^{\mathrm{in}}$ such that $g(W(x))$ is strictly positive and $\overline{g\circ W_{X^{\mathrm{in}}}}$ is strictly positive and finite. Finally, $(g\circ W)/\overline{g\circ W_{X^{\mathrm{in}}}}$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.

### v4

@@ 4.2 Summary of the $\mathcal{KL}_{\rm gen}$ extension
We extend the class of $\mathcal{K}\mathcal{L}$ functions that classically serve for dynamical systems stability. In our context, our $\mathcal{KL}_{\rm gen}$ class constitutes the main part of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound of $\varphi(T^{k}(x))$. We prove that such upper bounds exist if and only if $\nu$ has a finite supremum. Moreover, if ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, we can construct a function $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$ such that $e^{-1}\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$.
@@
Unfortunately, in practice, proving that a pair $(\gamma,\theta)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ seems to be as difficult as proving that a pair $(h,\beta)$ belongs to $\mathbb{\Gamma}(\nu)$. Thus, we propose a new type of Lyapunov function for which the conditions do not involve the entire trajectory of the analyzed dynamical system.
@@ 5 Lyapunov type Functions for Discrete-time Systems Maximization
Classically, Lyapunov functions (i.e., satisfying conditions (12) and (13)) are used as certificates of stability (see  [Ela05] or [KP01]). They present a constructive approach to prove the stability of a dynamical system through converse Lyapunov theorems, which prove that stability notions are equivalent to the existence of Lyapunov functions [Kha02]. For some linear, polynomial, or piecewise polynomial systems, Lyapunov functions can be computed from numerical optimization solvers based on linear or semidefinite programming (for example, see [GH15] and references therein). This numerical opportunity has been exploited for specific peak computation problems. Indeed, in [Adj25c] quadratic Lyapunov functions were used to solve our problem when the dynamical system is stable and affine and the objective function is quadratic. Next, in [Adj25a], we have proved that more general (local) Lyapunov functions can be used to solve peak computation problems. In these previous works, we applied direct-like Lyapunov theorems. In Lyapunov stability theory, after obtaining direct theorems, we consider converse theorems. The question raised in this case is, which stability notions imply the existence of a Lyapunov function? In our context, the converse theorem relies on the following question: which properties of classical Lyapunov functions can be removed to obtain an equivalence between the existence of a useful pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ and the existence of a weaker version of the Lyapunov function? In this section, we reply to this answer. First, in Subsection 5.1, we define a more general class of Lyapunov functions called Opt-Lyapunov functions. The existence of these functions does not ensure the stability of the system, but only the finiteness of the limit superior of the sequence $\nu$ defined in Eq.  (1). Then, in Subsection 5.2, we discuss the notion of usefulness and how it constrains the Opt-Lyapunov functions. In the classical case, usefulness constrains the sequence $\nu$ to have some positive terms. We replace the functions $\alpha_{1},\alpha_{\varphi}$ in (16) by a certificate of compatibility. Finally, in Subsection 5.3, we establish an equivalence theorem between the existence of those Opt-Lyapunov functions and the existence of an element $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
5.1 Generalized Lyapunov Functions for Discrete-time Systems Maximization
@@ 5.1.1 Definitions and Useful Facts
We extend the class of classical Lyapunov functions as follows: An Opt-Lyapunov function is still nonnegative but can be either strictly positive or can have several zeros and can take infinite values. The supremum of an Opt-Lyapunov function on the initial conditions set is non-null but smaller than 1. Opt-Lyapunov functions keep the decrease condition of classical Lyapunov functions.
@@
A function $V$ is said to be an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ if and only if
@@
Example 4 (Opt-Lyapunov function for the Leslie Model).
Returning to the Leslie model example introduced in Section 3. As $\lambda_{2}>1$, the system is unstable, that is, $\|x_{k}\|$ goes to $\infty$ as $k$ tends to $\infty$. A classical Lyapunov function does not exist for this system. However, an Opt-Lyapunov function exists. We propose, for $C>0$,
$\mathbb{R}^{2}\ni x\mapsto V_{\ell}(x)=\left\{\begin{array}[]{cr}\dfrac{C}{x_{1}^{2}+x_{2}^{2}}&\text{ if }x_{1},x_{2}>0\\ +\infty&\text{ otherwise}\end{array}\right.$. (32)
Let us take $x\in\mathbb{R}^{2}$. If $Lx$ has a nonpositive coordinate (and thus $V_{\ell}(Lx)=+\infty$), then either $x_{1}\leq 0$ or $x_{2}\leq 0$ yielding to $V_{\ell}(x)=+\infty$. Therefore, $V_{\ell}(Lx)\leq\lambda V_{\ell}(x)$ for any $\lambda\in(0,1)$. Now, assume that both coordinates of $Lx$ are strictly positive. This implies that $x_{1}>0$. We can have $x_{2}\leq 0$. In this case, $V_{\ell}(x)=+\infty$ and $V_{\ell}(Lx)\leq\lambda V_{\ell}(x)$ for any $\lambda\in(0,1)$. Finally, suppose that $Lx$ and $x$ have strictly positive coordinates. We have from  (20)
$\dfrac{V_{\ell}(Lx)}{V_{\ell}(x)}=\dfrac{x_{1}^{2}+x_{2}^{2}}{(ax_{1}+bx_{2})^{2}+cx_{1}^{2}}=\dfrac{x_{1}^{2}+x_{2}^{2}}{(a^{2}+c^{2})x_{1}^{2}+b^{2}x_{2}^{2}+2abx_{1}x_{2}}<\dfrac{x_{1}^{2}+x_{2}^{2}}{(a^{2}+c^{2})x_{1}^{2}+b^{2}x_{2}^{2}}<1.$
This inequality is not sufficient to prove condition (3) of Definition 4 as the supremum of $V_{\ell}(Lx)/V_{\ell}(x)$ might be equal to 1. We will compute the exact value of the supremum of $V_{\ell}(Lx)/V_{\ell}(x)$ over the open positive orthant of $\mathbb{R}^{2}$. This supremum is strictly smaller than 1.
We can adjust the constant $C$ to have $\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}}\in(0,1]$ even if we will see that the finiteness of $\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}}$ is sufficient to obtain the second statement of Definition 4. As $V_{\ell}$ is marginally strictly decreasing we have $\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}}=C/(l_{1}^{2}+l_{2}^{2})$. Thus any $C\in(0,l_{1}^{2}+l_{2}^{2}]$ can be chosen.
Applying those computations to the numerical data of Subsection 3.4, we get for all $x\in(\mathbb{R}_{+}^{*})^{2}$,
$V_{\ell}(x)=125/x_{1}^{2}+x_{2}^{2}$
where we choose $C=l_{1}^{2}+l_{2}^{2}$ to obtain $\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}}=1$. ∎
@@
If $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ then
@@
2. Let $x$ be a fixed point of $T$. If $V(x)=+\infty$, there is nothing to prove. Suppose that $V(x)$ is finite. From the third statement, $0\leq V(T(x))=V(x)\leq\lambda V(x)$ for some $\lambda\in(0,1)$. Hence, $V(x)(1-\lambda)=0$ and we conclude that $V(x)=0$. ∎
@@
If $X^{\mathrm{in}}$ is included in the fixed point set of $T$ then a function $V:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that $V\circ T\leq\lambda V$ for some $\lambda\in(0,1)$ cannot satisfy $\overline{V_{X^{\mathrm{in}}}}>0$.
Corollary 2 extends the fact that if a classical Lyapunov function exists for a self-map $T$, $T$ cannot have a nonzero fixed point. If an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ exists, $X^{\mathrm{in}}$ contains a vector $x$ such that $x\neq f(x)$.
In fact, we can equivalently replace the second statement of Def. 4 by $\overline{V_{X^{\mathrm{in}}}}$ is finite and strictly positive. Indeed, under this condition, we recover an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ with $\overline{W_{X^{\mathrm{in}}}}\in(0,1]$ by taking $W:=V/\overline{V_{X^{\mathrm{in}}}}$.
@@
The existence of a function $W:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that
@@
is equivalent to the existence of an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
@@
Let $\gamma$ be strictly greater than 1. Let $f\in\mathcal{K}_{\infty}$ verifying for all $x>0$, $f(x)>x$. Then, there exists a function $g\in\mathcal{K}$ that satisfies $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$.
Lemma 3 is a slightly modified version of [Kel14, Lemma 25], for which $f-\mathrm{Id}$ is supposed to be in $\mathcal{K}_{\infty}$. To be self-contained, we propose in the Appendix a proof inspired by the one provided in [Kel14].
@@
Again, a proof of Lemma 4 is given in [Kel14]. A proof is provided in the Appendix for the convenience of the reader.
Proof of Prop. 3.
Suppose that such $W$ and $\rho$ exist, then let $\lambda\in(0,1)$. From Lemma 4, there exists $g$ such that $g(\rho(x))=\lambda g(x)$ for all $x\in\mathbb{R}_{+}$. By extension, we define $g(+\infty)=+\infty$. Then as $g$ is increasing, we have $g(W(T(x)))\leq g(\rho(W(x))=\lambda g(W(x))$ for all $x\in\mathbb{R}^{d}$. Now, as $g$ is strictly increasing, there exists some $x\in X^{\mathrm{in}}$ such that $g(W(x))$ is strictly positive and $\overline{g\circ W_{X^{\mathrm{in}}}}$ is strictly positive and finite. Finally, $(g\circ W)/\overline{g\circ W_{X^{\mathrm{in}}}}$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
@@
Prop. 4 shows that the set of Opt-Lyapunov functions contains the set of classical Lyapunov functions.
Proposition 4 (Classical is Opt).
If $V$ is a classical Lyapunov function for $T$ and $X^{\mathrm{in}}$ is bounded and not reduced to $\{0\}$, then $V/\overline{V_{X^{\mathrm{in}}}}$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
Proof.
Let $V$ be a classical Lyapunov function for $T$. If $X^{\mathrm{in}}$ is not reduced to $\{0\}$, as $V$ is positive definite, $V(x)>0$ for some $x\in X^{\mathrm{in}}$. From  (12), if $X^{\mathrm{in}}$ is bounded, $\overline{V_{X^{\mathrm{in}}}}$ is finite. This implies that $V/\overline{V_{X^{\mathrm{in}}}}$ satisfies the second statement in Definition 4. Moreover, it is obvious that $V/\overline{V_{X^{\mathrm{in}}}}$ is positive definite and so verifies the first statement of Definition 4. It is easy to see that $V/\overline{V_{X^{\mathrm{in}}}}$ satisfies the third statement of Definition 4 from (13) and $\overline{V_{X^{\mathrm{in}}}}>0$. ∎